Help me please:
6n= 2/3(5n-2)

Answer by rothauserc(4718)

6n= 2/3(5n-2)
multiply both sides of = by 3
18n = 2*(5n-2)
expand right side of =
18n = 10n-4
subtract 10n from both sides of =
8n = -4
n = -4/8 = -1/2